dernesha:

What would be the value of $150 after eight years if you earn 12 percent interest per year?

jhonyy9:

12% interest per year of $150 using the rule of 3 simple calcule how many is the 12% of $150 12 from 100 x from 150 so x = 12*150/100 = 18 so this is the interst rate per one year but the question ask per 8 years so just you need multiplie the 18 by 8 18×8 = 144 so this is the interest rate per 8 years assuming 150 with 144 you get the right answer what will be 150+144 = 294 so the value of $150 after 8 years will be $294 hope this is understandably and hope helped